OpenStudy (anonymous):

Find x in this 45-45-90 degree triangle. x=leg hypotenuse=sqrt of 24 (put into radical form- number times sqrt of another number)

OpenStudy (anonymous):

x=2 times square root of 3

OpenStudy (anonymous):

a^2 + b^2 = c^2 so we have c^2 = to sqrt(24) now if we re-write it, it will look like this x^2 + x^2 = (sqrt(24))^2 then all that is left is some algebra, does this help?

OpenStudy (anonymous):

It helps for the moment while I remember it.

OpenStudy (anonymous):

with terms that have the same base and exponents, we can re-write them like this x^2 + x^2 becomes 2x*2 because they have the same base and the same exponent and we are adding them together so we then would have 2x^2 = (sqrt(24))^2